# Discrete Diffusion VLA: Bringing Discrete Diffusion to Action Decoding in Vision-Language-Action Policies

### v1

@@ 4.3 Ablation Study
Decoding strategy. We compare one-shot parallel decoding, random order, confidence-gap selection, and our max-confidence selection. On LIBERO-Goal the success rates are 95.6%, 96.0%, 96.6%, and 97.4% respectively. The adaptive easy-first schedule guided by per-token confidence gives +1.8 points over one-shot parallel and +0.8 over confidence-gap, while random order lags behind. These results indicate that ranking tokens by instance-wise confidence and resolving the easy elements first improves refinement efficiency and final accuracy.
Choice temperature. We study the temperature used to turn posteriors into discrete choices during refinement. Hard argmax at all steps reaches 96.2%, a fixed temperature of 1.0 gives 96.4%, and a simple linear decay from 1.0 to 0.0 across steps attains 97.4% on LIBERO-Goal. The decayed schedule encourages mild exploration early and sharper commitment later, and it works well with adaptive decoding to correct early mistakes and consolidate consistent actions.
Table 4: Ablation study. Success rates on LIBERO-Goal when varying the decoding strategy and the choice temperature.
Ablation Item Decoding Strategy Choice Temperature
@@
[…]
4.4 Analysis on Inference Efficiency
A key advantage of our discrete diffusion approach lies in its inference efficiency, particularly regarding NFEs compared to AR models. For an action chunk of length $L=H\times D_{\text{act}}$, an AR decoder without mechanisms like Multi-Token Prediction (Gloeckle et al., 2024) must perform $L$ sequential forward passes, as each token prediction is conditioned on all previously generated ones. This results in an NFE count of exactly $L$. For instance, in our LIBERO experiments where the horizon $H=8$ and action dimension $D_{\text{act}}=7$, the total chunk length is $L=56$. An AR model would thus require 56 sequential function evaluations, creating a computational bottleneck where latency scales linearly with the prediction horizon. It’s a critical constraint for real-time robotics applications.
In contrast, our Discrete Diffusion VLA denoises the entire action chunk over a small, fixed number of refinement steps, $T$. At each step $t$, the model performs a single forward pass to predict posteriors for all currently masked tokens simultaneously, making the total NFE equal to $T$. In our experiments, we use a default of $T=12$ steps. This reduces the required NFEs from $L=56$ to a constant $T=12$, a 4.7x reduction in function evaluations for the LIBERO setup. Discrete diffusion decouples the inference cost from the sequence length. The adaptive decoding and secondary re-masking mechanisms are lightweight computations performed on the output logits and do not require additional forward passes, preserving the core efficiency of the parallel refinement process.
@@ 5 Conclusion
Discrete Diffusion VLA unifies vision, language, and action inside a single transformer by marrying diffusion’s progressive-refinement paradigm with a discrete-token action interface, enabling an adaptive “easy-first, hard-later” decoding order and secondary re-masking for reliable error correction. Our architecture preserves pretrained VLM priors, the left-to-right bottleneck common in autoregressive VLA, and delivers state-of-the-art performance across LIBERO and two SimplerEnv suites while using fewer function evaluations than continuous-diffusion and autoregressive baselines. By aligning action decoding with the VLM transformer, Discrete Diffusion VLA offers a path to inherit unified-transformer scaling behavior, paving the way for large-scale VLA with larger models and broader datasets.
Limitations: Our fixed-bin action tokenization remains coarse relative to continuous control, leading to sub-bin precision loss. Developing more expressive action encoding, or integrating a mechanism that imposes the continuous-space loss with discrete diffusion VLA, is a promising direction.
@@ Acknowledgments
We sincerely thank Wanqi Zhong from Harbin Institute of Technology for assistance with figure illustration and layout.

### v2

@@ 4.3 Ablation Study
Decoding strategy. We compare one-shot parallel decoding, random order, confidence-gap selection, max-confidence selection and our max-confidence + secondary remasking. On LIBERO-Goal, the success rates are 95.6%, 96.0%, 96.6%, 97.0%, and 97.4% respectively (Tab. 4). The adaptive easy-first schedule guided by per-token confidence yields $1\%$ over one-shot parallel. And max-confidence selection is slightly better than confidence-gap, while random order lags behind. Additionally, our secondary remasking scheme also brings about $0.4\%$ benefits. These results indicate that ranking tokens by instance-wise confidence and resolving the easy elements first improves refinement effectiveness and final accuracy.
Choice temperature. We study the temperature used to turn posteriors into discrete choices during refinement. Hard argmax at all steps reaches 96.2%, a fixed temperature of 1.0 gives 96.4%, and a simple linear decay from 1.0 to 0.0 across steps achieves 97.4% (Tab. 5). Decay encourages mild exploration early and sharper commitment later, complementing adaptive decoding to correct early mistakes and consolidate consistent actions.
Denoise Steps. Please refer to Fig. 4. And detailed illustrations are in the following section.
Table 4: Ablation study on decoding strategy. (LIBERO-Goal). Ranking tokens by instance-wise confidence improves over one-shot parallel by >1>1 point, and our max conf+secondary remask yields the best accuracy (97.4%).
Decoding Strategy Parallel Decoding Random Order Confidence Gap Max Confidence Max Conf+ Second Remask
Success 95.6% 96.0% 96.6% 97.0% 97.4%
Table 5: Ablation study on choice temperature. (LIBERO-Goal). Linear decay temperature attains the best 97.4%, encouraging mild exploration early and sharper commitment later.
Choice Temperature Hard Sample (Temp=0) Fixed Temp (Temp=1) Linear Decay Temp (Temp=1-$t$)
Success Rates 96.2% 96.4% 97.4%
Figure 4: Speed–Quality trade-off. (i) Time efficiency by the number of generated action chunks per second. (ii) Ablation on denoising steps. We adopt $T=12$ as a knee point for high accuracy at strong throughput.
4.4 Analysis on Inference Efficiency
Another benefit of discrete diffusion is inference efficiency in number of function evaluations (NFEs) compared to autoregressive decoding. For an action chunk of length $L{=}H{\times}D_{\text{act}}$, an AR decoder without mechanisms like multi-token prediction (Gloeckle et al., 2024) performs $L$ sequential forward passes (each token conditions on all previous ones). For instance, in LIBERO, H=8H{=}8 and $D_{\text{act}}{=}7$ yield L=56L{=}56, creating a latency bottleneck that scales linearly with horizon. It’s a critical constraint for real-time robotic applications.
@@ 4.4 Analysis on Inference Efficiency
Speed–quality trade-off. Figure 4 sweeps $T$ and reports throughput (number of action chunks generated per second; left $y$-axis) and success rates on LIBERO-GOAL (right $y$-axis). Accuracy improves monotonically with $T$ but with diminishing returns: most gains appear by $T{\in}[8,12]$, while increasing to $T{\geq}16$ yields sub–1 point improvements at a disproportionate drop in speed (throughput scales roughly as $1/T$). We therefore choose T=12T{=}12 as a knee point that matches the best accuracy in Tables 45 while maintaining high throughput (about 3 Hz for action chunk generation).
@@ 5 Conclusion
Discrete Diffusion VLA unifies vision, language, and actions inside a single transformer by marrying diffusion’s progressive-refinement paradigm with a discrete-token action interface, enabling an adaptive “easy-first, hard-later” decoding order and secondary re-masking for reliable error correction. Our architecture preserves pretrained VLM priors, breaks the left-to-right bottleneck in autoregressive VLA, and delivers state-of-the-art performance across LIBERO and two SimplerEnv suites while using fewer function evaluations than AR baselines. By aligning action decoding with the VLM transformer, Discrete Diffusion VLA offers a path to inherit unified-transformer scaling behavior, paving the way for large-scale VLA research with larger models and broader datasets.
@@ Acknowledgments
We sincerely thank Wanqi Zhong from Harbin Institute of Technology for assistance with figure illustration and layout. This project is partially supported by Chinese Institute of Electronics-Tencent Doctoral Research Incentive Program.
Appendix A Disclosure of Large Language Model Usage
We employed large language models (LLMs) only for minor copyediting—grammar and phrasing. All LLM-suggested edits were carefully reviewed and verified by the authors to prevent fabricated content and preserve the original intent. The research ideas, methodology, implementation, experiments, data analysis, and conclusions presented in this work were entirely conceived and executed entirely by the authors without LLM assistance.
Appendix B Visualizations of Robot Task Executions
i) Franka Panda Arm on LIBERO-Spatial Task
ii) Google Robot on Move Near Task
iii) WidowX Arm on Put Eggplant in Basket Task
iv) Franka Panda Arm on LIBERO-Long Task
v) WidowX Arm on Stack Green Block on Yellow Block Task
Appendix C Implementation Details
  1. 1.
    We choose action chunk size $H$ as $8$ for both LIBERO and SimplerEnv-Fractal, and $3$ for SimplerEnv-Bridge, following widely used settings in each environment, respectively.
  2. 2.
    We conduct our experiments with batch size $32$ for all of the experiments and typically conduct each run on 4 NVIDIA A800 TENSOR CORE GPUs.
  3. 3.
    We apply Feature-wise Linear Modulation (FiLM) (Kim et al., 2025b) on Simpler-Bridge experiments to enhance the language grounding abilities of our model on WidowX Arm manipulation tasks.
  4. 4.
    We train our model on LIBERO-Spatial and LIBERO-Object for 150k steps while 300k steps on LIBERO-Goal and LIBERO-Long. And we report the results of each highest checkpoints. Besides, we only train 100k steps on both Simpler-Fractal and Simpler-Bridge and report the highest overall performance of each environment.
  5. 5.
    For secondary re-masking, we set $\eta^{abs}_{t}=0.5\times(1-t/T)$ as the step-dependent threshold of threshold check.